The principal amount of the loan borrowed is $12,000, origination fee of $50, interest rate is 6.20%. With no payments due before, what will the payoff be in 6 months?
Answer #1
Calculation of Loan payoff in 6 months
We can calculate the loan payoff using future value of sum formula stated under,
FV = P * (1+r)^n
FV = Loan payoff = ?
P = principal amount of loan borrowed = $12000
r = interest rate per month = 6.20%/12 = 0.005167
n = number of months = 6
FV = 12000 * (1+0.005167)^6
FV = 12000 * 1.03140
FV = 12376.84
Loan payoff in 6 months = $12,376.84
Note
An origination fee is an upfront fee charged by a lender for processing a new loan application.
Hence it is not considered in calculation of payoff.